The electric flux f through a hemisphere surface of radius R, placed in a uniform electric field of intensity


E parallel to the axis of its circular plane is:


(a) 2 p RE (b) 2 pR2E (c) pR2E (d) (4/3) pR3E


 


Grade:11

3 Answers

Anirban Mukherjee
11 years ago

Electric Flux through any surface

= Electric field Intensity at the surface

´ component of surface ^

to direction of electric Field

=EΠR2

Sharandeep singh
11 years ago

since it is placed in uniform e field therefore net charge enclosed is 0 and therefore net flux is zero

i.e. flux through circular part + flux through hemi spherical part = 0

flux through circular part = - flux through hemi spherical part

                                   = e * 2pr^2 ( area of hemisphere = 1/2*4pr^2 = 2pR^2 )

hence b is correct

Kushagra Madhukar
4 years ago
Dear student,
Please find the solution to your question below.
 
Electric Flux through any surface = Electric field Intensity at the surface x component of surface perpendicular to direction of electric Field = EΠR2
 
Hope it helps.
Thanks and regards,
Kushagra